OpenStudy (anonymous):

Suppose that 2 ≤ f '(x) ≤ 5 for all values of x. What are the minimum and maximum possible values of f(6) − f(3)? ATTEMPT: Unfortunately for this one, I am not sure how I am going to proceed with it. If anyone cares to help or give a hint, that would be great. The answer comes in blanks like this: [ ] ≤ f(6) − f(3) ≤ [ ] Suppose that 2 ≤ f '(x) ≤ 5 for all values of x. What are the minimum and maximum possible values of f(6) − f(3)? ATTEMPT: Unfortunately for this one, I am not sure how I am going to proceed with it. If anyone cares to help or give a hint, that would be great. The answer comes in blanks like this: [ ] ≤ f(6) − f(3) ≤ [ ] @Mathematics

OpenStudy (anonymous):

Wow how did you get to this page so fast

OpenStudy (anonymous):

the calculus call to me

OpenStudy (anonymous):

HAHA

OpenStudy (cathyangs):

It would be 0\[0\le f(6)-f(3) \le 3\] I believe

OpenStudy (anonymous):

Notice that f is always increasing

OpenStudy (anonymous):

Aw shoot. I should've said I didn't want an answer. Its alright though. The reason being is that I don't really understand if I get the answer in front of me. I can get that sensation by looking at the back of the textbook.

OpenStudy (anonymous):

Yes f is increasing throughout the entire interval

OpenStudy (cathyangs):

oh... I'm sorry.

OpenStudy (anonymous):

another hint: secant line

OpenStudy (anonymous):

Oh I see, so the end values of the interval I need to use them?

OpenStudy (cathyangs):

wait...but it doesn't say that f(n) increases as n increase. I figured that the smallest difference would be 0, because two n can produce the same f(n) result. And the biggest would be 3, where one is 2 and the other is 3.

OpenStudy (cathyangs):

ah!!!! I gave you the wrong answer way up there ^, not the previous reply, but the first one that I put.

OpenStudy (anonymous):

Wait a second. Can I use antiderivation to solve this?

OpenStudy (anonymous):

yes, but if we notice that f' is between 2 and 5 it is increasing in that interval, which denote that the slope is positive, which further means that the orginal function is increasing

OpenStudy (anonymous):

So what if the function is increasing?

OpenStudy (anonymous):

it help us because now we are able to use the secant line to determine or solve this inequality, since it will have a slope between 2 and 5

OpenStudy (jamesj):

Equivalently, by the Fundamental Theorem of Calculus, \[ f(6) - f(3) = \int_3^6 f'(x) \ dx \] Now as \( 2 \leq f'(x) \leq 5 \), it follows that \[ \int_3^6 2 \ dx \leq \int_3^6 f'(x) \ dx \leq \int_3^6 5 \ dx \] Hence ... I'm sure you've got it from here.

OpenStudy (jamesj):

got it?

OpenStudy (anonymous):

I'll work on it from here. Thanks for your help.

OpenStudy (jamesj):

tell me what final answer you get. It's actually a nice problem.